# Sphere Packing on a Quantum Computer for Chromatography Modeling

### v1

@@ II.3.3 Hyperparameter Optimization
The Lagrange multiplier $\lambda$ in the integer program for MIS (Eq. 4) must be chosen carefully. If $\lambda$ is too small, the energies associated with invalid packing will not be sufficiently separated from valid packings. This will lead to QAOA including infeasible states in its search, thus hampering both its search efficiency and accuracy. However, if $\lambda$ is too large, it will result in wild swings in energy for the classical minimization algorithm. This hampers the ability of the classical computer to minimize the cost landscape. To find a $\lambda$ that achieves the “sweet spot” between these two extremes, we can treat $\lambda$ as a hyperparameter, which can be optimized via a hyper-loop. To objectively judge how well a given $\lambda$ helps QAOA, we plot the probability that QAOA will return the optimal packing vs the number of layers $p$ for various $\lambda$ (Figure 9). The plot shows that $\lambda=0.5$ yields the highest probability of finding the optimal state as $p$ grows while additionally finding the optimal state with the highest probability. Thus $\lambda=0.5$ is chosen for running QAOA on real quantum hardware.
Given the increasing optimal state overlap with decreasing $\lambda$, one may wonder why we don’t select and even smaller $\lambda$. However, upon further numerical simulations, we saw that for smaller $\lambda$’s, namely $\lambda=0.25$, the energy penalty for invalid packings was so small that, even after 20,000 runs of a noiseless simulation, QAOA couldn’t find the optimal solution even once (across all $p$). This would constitute and example of $\lambda$ being choose to be too small, as the small energy penalty allowed invalid packings to swamp the optimizer and preclude it from finding the optimal solution.
Figure 9: Probability of optimal state versus number of layers for various $\lambda$. A circle ($\bullet$) or cross ($\times$) indicates that the optimal state was or wasn’t the most probable state measured, respectively.
@@ II.3.4 Compilation
Plugging the MIS Hamiltonian (Eq. 4) into the QAOA ansatz (Eq. 1) results in a set of $R_{ZZ}$, $R_{Z}$, and $R_{X}$ gates. The $R_{ZZ}$ operation can be decomposed into single-qubit gates and exactly two $CZ$ gates (Garnet’s native two-qubit gate). Meanwhile, to reduce total gate count, the $R_{Z}$ and $R_{X}$ gates can be combined (as $R_{Z}$ commutes through $R_{ZZ}$) and further expressed in terms of $PR_{X}$ gates (phased-$R_{X}$ gates, $IQM^{\prime}s$ native single qubit gate). Furthermore, in an effort to minimize the depth of the quantum circuit, we partition the $R_{ZZ}$ gates into 4 disjoint sets, corresponding to the maximum edge coloring of the coupling map (Figure 10).
Figure 10: Maximum edge coloring of the Garnet coupling map.
@@ II.3.5 Noisy Simulation
Before executing our QAOA circuits on hardware, we considered an ideal simulation of the QAOA circuit with $\lambda=0.5$ over increasing circuit depth ($p$), and found, as expected, that QAOA can find the optimal solution with greater probability as the depth increases (Figure 11). However, due to the Noisy Intermediate-Scale Quantum (NISQ) [13] stage of current devices, the existence of a consistent level of uncorrected noise in certain systems, obtaining good results can be hampered as circuit gate counts increase. That is, there is a balance to be had (at least for NISQ devices) in the circuit depth that can be run before the increasing noise from gate count overwhelms QAOA from finding the optimal solution with the highest probability. To that end, we additionally ran a noisy simulation of our QAOA circuits to predict which depth $p$ would perform best on the selected hardware (tailoring our noise parameters to IQM’s Garnet QPU).
Figure 11: Plot of an ideal and noisy simulation of compiled QAOA circuits with increasing depth, $p$. The simulations were done by sampling 20,000 shots, using pre-computed angles, and $\lambda=0.5$. The noisy simulation was subject to single-, two-qubit, and readout errors (Subsection II.3.5) modeled specifically for IQM’s Garnet QPU.
@@
We started by optimally compiling the circuit, matching the qubit connectivity, and thus reducing the need to consider error sources arising from circuit routing and its resultant introduction of noisy SWAP gates. We then modeled single- and two-qubit errors in addition to measurement readout errors. The device noise calibration data was obtained from AWS Braket during the time of the experiment. The single-qubit errors consisted of an amplitude damping channel based on the $T_{1}$ thermal relaxation times reported for each single qubit, then a phase damping channel on each qubit following its respective $T_{2}$ de-phasing time. Lastly, we included a single-qubit depolarizing channel with the depolarization rate taken to be roughly as one minus the fidelity of each respective qubit’s simultaneous randomized benchmarking result. Given that the only two-qubit gate in our QAOA circuits is the $CZ$ gate, we employed a two-qubit depolarizing noise channel to model that gate error source – with the depolarization rate taken to be the reported $CZ$ infidelity per qubit pair in the coupling map (applied for both gate directions). Finally, we modeled the expected readout error of measurements on hardware via a bit flip channel with a probability of one minus the respective qubit’s readout rate. Combining all these channels, and applying them per gate operation, we plotted how the probability of the optimal solution varies as a function of $p$ in Figure 11. Following our earlier expectation, we find that the various noise sources lower our probabilities and we also understandably find that, after $p=3$, the probability of the optimal solution actually decreases – contrary to the ideal case scenario.
@@
This highlights that the limitation stems more from the noise levels of current hardware than from the ability of QAOA to perform effectively. We therefore expect that quantum computers becoming less noisy, along with a combination of quantum error mitigation, suppression, and correction, will help overcome this discrepancy. In the current term though, carrying out a noisy simulation can help inform the circuit depths that could be optimally run on hardware before qubit noise starts to become a significantly diminishing factor, as we have done here.
@@ II.3.6 Experiment Results
Figure 12: Probabilities of the top 10 most probable states measured on Garnet for $p=1,3,5$. The opaque bars represent the optimal solution while the transparent bars represent non-optimal solutions.
To execute our circuits on Garnet, we compiled and submitted our jobs through Infleqtion’s compilation software, Superstaq [14] which, in turn, submitted the job through AWS’s Braket. We ran QAOA for $p=3$ (based on the noisy simulations of Subsection II.3.5) and also $p=1$ and $p=5$ for comparison. Each was executed with 20,000 shots (the maximum number of shots allowed by IQM for submissions to Garnet through AWS). The variational parameters $\alpha$ and $\beta$ were trained via classical simulation and the final quantum circuit was run on the quantum device using these optimized parameters. In the small case of this example experiment, the classical computer can simulate the quantum circuit; however, at larger problem sizes, one may take advantage of parameter concentration (Section II.3.7) to train the variational parameters via classical simulation on a smaller sub-problem and then run the quantum circuit for the full problem with these optimized parameters on a quantum computer. At a large enough scale, the quantum computer becomes necessary for the running of the final circuit as classical simulation of quantum circuits scales exponentially with the number of qubits.
Figure 12 shows the probabilities of the top ten most probable states measured on Garnet for each $p$. The opaque bars represent the optimal solution while the translucent bars represent non-optimal solutions. Note that for each $p$, the most probable state was the optimal state. The figure also demonstrates the two competing forces involved when selecting the best $p$. On the one hand, increasing $p$ increases the expressiveness of the circuit and therefore the effectiveness of QAOA. On the other hand, increasing $p$ also increases the depth of the quantum circuit, and therefore the noisiness of the results. For example, the plot for $p=1$ is sharp because of its short depth, yet low probability because of its low expressivity. By contrast, the plot for $p=3$ is high probability because of its high expressivity, yet flat because of its long depth. The ‘sweet spot’ is at $p=3$ (as also suggested by our noisy simulations from Subsection II.3.5), where the circuit is expressive enough to render high probabilities yet short enough to be sufficiently sharp. As one expands to larger, more expensive-to-run problem sizes, the importance of preemptively searching for the “sweet spot” for $p$ via noisy simulations of the circuit before running on hardware (Subsection II.3.5) becomes ever more useful.
@@ II.3.7 Parameter Concentration
Figure 13:
Figure 14:
As noise often scales with the number of qubits/gates, and therefore problem size, it becomes harder to train QAOA as the problem size increases. This roadblock can be alleviated, however, if the problem exhibits parameter concentration [15], a phenomenon in which the QAOA parameters trained on a smaller instance of a combinatorial optimization problem perform well when applied to a larger instance of the same problem. Fortunately, we have numerical evidence suggesting parameter concentration for our particular problem of sphere packing. Figure 13 shows the full graph used for the hardware experiment. The dots colored red represent a subgraph which will serve as the “smaller problem instance” for our parameter concentration investigation. We train the parameters of both the subgraph QAOA circuit ($C_{\text{sub}}$) and the full graph QAOA circuit ($C_{\text{sub}}$), resulting in $\rho_{\text{sub}}$ and $\rho_{\text{full}}$, respectively. To test for parameter concentration, we compare the performance of the full graph circuit executed with the parameters trained on the subgraph, $C_{\text{full}}(\rho_{\text{sub}})$, with that of the full graph circuit executed with the parameters trained on the full graph, $C_{\text{full}}(\rho_{\text{full}})$. Figure 14 compares the performance of $C_{\text{sub}}(\rho_{\text{sub}})$, $C_{\text{full}}(\rho_{\text{sub}})$, and $C_{\text{full}}(\rho_{\text{full}})$ with respect to overlap with the optimal state for various numbers of layers $p$, respectively. We can see that QAOA yields a larger overlap with the optimal state when solving the subgraph with the parameters trained on the subgraph $C_{\text{sub}}(\rho_{\text{sub}})$ when compared to solving the full graph with the parameters trained on the full graph $C_{\text{sub}}(\rho_{\text{sub}})$. This is expected as the subgraph has less parameters and is therefore easier to train. However, we can see that QAOA on the full graph with parameters trained on the subgraph $C_{\text{full}}(\rho_{\text{sub}})$ performs only slightly worse than QAOA on the full graph with parameters trained on the full graph $C_{\text{full}}(\rho_{\text{full}})$, suggesting parameter concentration. This suggests that QAOA may be able to solve real-world instances of this problem that are too large to train by executing QAOA on parameters trained on a trainable sub-instance.
@@ III Heterogeneous Packing
Using the encouraging results from real quantum hardware as a springboard, we generalize sphere packing to the heterogeneous case, for which good classical solutions are much harder to achieve. Consider the case of $s+1$ distinct radii $R=\{r_{n}\}_{n\in N}$ where $N=\{0,1,...,s\}$ and $r_{0}=0$ (a useful indicator for the nonexistence of a sphere). The state of the qubit (or set of qubits) at each node informs which sized sphere is centered at said node. To maximize the tightness of the packing, we minimize empty space subject to the constraint that the arrangement of spheres is non-overlapping. We formulate the cost Hamiltonian $H_{C}=H_{obj}+\lambda H_{con}$ for heterogeneous packing via two different formulations, which we call “First Quantization” and “Second Quantization” because of their analogies to the two levels of quantization in physics.
@@
$N_{vw}$ =$=\{n,m\in N_{v}\times N_{w}:|v-w|\leq r_{n}+r_{m}\}$ (9)
$N_{v}$ represents the index set of the set of radii that admit spheres of said radii to be placed at node $v$ without extending beyond the boundary of the cylinder. Similarly, $N_{vw}$ represents the index set of pairs of radii that would not admit the placement of spheres of said radii at nodes $v$ and $w$ without them overlapping. Here, $R_{b}$ is the radius of the cylinder. Converting these Hamiltonians to a qubit formulation (a requirement for any real hardware system today) requires only $\log_{2}|R|$ overhead via a binary encoding.
@@ IV Resource Scaling
Resource estimation for this problem is important to argue for the potential for quantum advantage over classical computation. To estimate this scaling of quantum resources for each problem formulation, we start by making several reasonable assumptions. Suppose we have a $d$-dimensional lattice with $q$ points per side placed on a $d$-dimensional cylinder with radius $R_{b}$. Furthermore, notate the set of radii, node sets, edge sets and the maximum radius as
@@
. The number of CNOT gates to implement $e^{iH\gamma}$ in one QAOA layer is.
$\sum_{r_{i},r_{j}\in R^{2}}|E_{r_{i}r_{j}}|$ $\leq|R|^{2}q^{d}\max_{r\in R}\left(\max_{v\in V_{r}}\deg(v)\right)$ (16)
$\leq|R|q^{d}\left(\frac{r_{m}q}{R_{b}}\right)^{d}$ (17)
@@
$R$ =$=\{r_{1},...,r_{s}\}$ (19)
$\Pi$ =$=\{1,...,\lceil\log_{2}\left(|R|+1\right)\rceil\}$ (20)
$V$ =$=\{v\}$ (21)
@@
where $E_{r_{i}r_{j}}$ has an edge $(v,w)$ if and only the distance between nodes $v$ and $w$ is less than $r_{i}+r_{j}$. With these definitions, the number of qubits is simply:
$|\Pi||V|$ $\leq\left(\lceil\log_{2}\left(|R|+1\right)\rceil\right)q^{d}$ (23)
@@
$g(v,r_{i},w,r_{j})=\begin{cases}1&\text{if }(v,w)\text{ }\\ 0&\text{else}\end{cases}$ (25)
@@
$\sum_{r_{i},r_{j}\in R^{2}}\sum_{v,w\in E_{r_{i}r_{j}}}$ $\leq|R|2^{2|\Pi|}\max_{r_{i},r_{j}\in R^{2}}\left(\max_{v\in V}\deg_{r_{i}r_{j}}(v)\right)$ (26)
$\leq|R|2^{\left({2\lceil\log_{2}\left(|R|+1\right\rceil}\right)}\left(\frac{r_{m}q}{R_{b}}\right)^{d}$ (27)
@@
Note that the bound on the single state penalty function is quite loose. Additionally, while the scaling may appear exponential in $d$, recall that for real-world applications $d$ is only $2$ or $3$, keeping the quantum resource scaling bounded by a low order polynomial in all its components.

### v2

@@ II.3.3 Hyperparameter Optimization
Figure 9: The probability of measuring the optimal state versus the number of layers for various $\lambda$. A circle ($\bullet$) or cross ($\times$) indicates that the optimal state was or wasn’t the most probable state measured, respectively.
The Lagrange multiplier $\lambda$ in the integer program for MIS (Eq. 4) must be chosen carefully. If $\lambda$ is too small, the energies associated with invalid packing will not be sufficiently separated from valid packings. This will lead to QAOA including too many infeasible states in its search, thus hampering both its search efficiency and accuracy. However, if $\lambda$ is too large, it will result in wild swings in energy for the classical minimization algorithm, hampering its ability to minimize the cost landscape. To find a $\lambda$ that achieves the “sweet spot” between these two extremes, we treat $\lambda$ as a hyperparameter, which is then optimized via a hyper-loop. To objectively judge how well a given $\lambda$ helps QAOA, we plot the probability that QAOA will return the optimal packing vs number of layers $p$ for various $\lambda$ (Figure 9). The plot shows that $\lambda=0.5$ yields the highest probability of finding the optimal state as $p$ grows while additionally finding the optimal state with the highest probability for all $p$. Thus $\lambda=0.5$ is chosen for running QAOA on real quantum hardware.
Given the increasing optimal state overlap with decreasing $\lambda$, one may wonder why we don’t select and even smaller $\lambda$. However, further numerical simulations indicated that for smaller $\lambda$’s, namely $\lambda=0.25$, the energy penalty for invalid packings was so small that, even after 20,000 runs of a noiseless simulation, QAOA couldn’t find the optimal solution even once (across all $p$). This would constitute an example of $\lambda$ being too small, as the small energy penalty allowed invalid packings to swamp the optimizer, precluding it from finding the optimal solution.
@@ II.3.4 Compilation
Plugging the MIS Hamiltonian (Eq. 4) into the QAOA ansatz (Eq. 1) results in a set of $\text{R}_{ZZ}$, $\text{R}_{Z}$, and $\text{R}_{X}$ gates. The $\text{R}_{ZZ}$ operation can be decomposed into single-qubit gates and exactly two CZ gates (Garnet’s native two-qubit gate). Meanwhile, to reduce total gate count, the $\text{R}_{Z}$ and $\text{R}_{X}$ gates can be combined (as $\text{R}_{Z}$ commutes through $\text{R}_{ZZ}$) and further expressed in terms of $\text{PR}_{X}$ gates (phased-$\text{R}_{X}$ gates, IQM’s native single qubit gate). Furthermore, in an effort to minimize the depth of the quantum circuit, we partition the $\text{R}_{ZZ}$ gates into 4 disjoint sets, corresponding to the maximum edge coloring of the coupling map (Figure 10). Because no edges of the same color share a node, every $\text{R}_{ZZ}$ corresponding to an edge of the same color may be applied simultaneously. Thus, all $\text{R}_{ZZ}$ gates may be applied with only depth 4, the smallest possible number of disjoint sets for this graph.
Figure 10: Maximum edge coloring of the Garnet coupling map. The $\text{R}_{ZZ}$ gates between edges of the same color are applied simultaneously.
@@ II.3.5 Noisy Simulation
Before executing our QAOA circuits on hardware, we considered an ideal simulation of the QAOA circuit with $\lambda=0.5$ over increasing circuit depth $p$ and found that, as expected, the probability of QAOA measuring the optimal solution increases with increasing depth (Figure 11). However, because we are currently in the Noisy Intermediate-Scale Quantum (NISQ) [20] era, the existence of a consistent level of uncorrected noise makes obtaining good results difficult as circuit depth increases. That is, there is a balance to be had (at least for NISQ devices) in the circuit depth that can be run before the increasing noise from gate count overwhelms QAOA from finding the optimal solution with the highest probability. To this end, we additionally ran a noisy simulation of our QAOA circuits to predict which depth $p$ would perform best on the selected hardware (tailoring our noise parameters to those of IQM’s Garnet).
@@
We started by optimally compiling the circuit (Subsection II.3.4) thus reducing the need to consider noise arising from circuit routing and its resultant introduction of noisy SWAP gates. We then modeled single- and two-qubit errors in addition to measurement readout errors. The device noise calibration data were obtained from AWS Braket shortly before the time of the experiment. The single-qubit errors consisted of an amplitude damping channel based on the $T_{1}$ thermal relaxation times reported for each single qubit, and a phase damping channel on each qubit following its respective $T_{2}$ de-phasing time. Lastly, we included a single-qubit depolarizing channel with the depolarization rate taken to be roughly as one minus the fidelity of each respective qubit’s simultaneous randomized benchmarking result. Given that the only two-qubit gate in our QAOA circuits is the CZ gate, we employed a two-qubit depolarizing noise channel to model that gate error source – with the depolarization rate taken to be the reported CZ infidelity per qubit pair in the coupling map. Finally, we modeled the expected readout error of measurements on hardware via a bit flip channel with a probability of one minus the respective qubit’s readout rate. Executing a noisy-simulation using all of these channels per gate operation resulted in the blue line of in Figure 11. Following our earlier expectation, we find that the various noise sources lower our probabilities and we also understandably find that, after $p=3$, the probability of the optimal solution actually decreases – contrary to the ideal case scenario.
Figure 11: Plot of an ideal and noisy simulation of compiled QAOA circuits with increasing depth $p$. The simulations were done by sampling 20,000 shots, using pre-computed angles, and setting $\lambda=0.5$. The noisy simulation was subject to single-qubit, two-qubit, and readout errors which model IQM’s Garnet.
@@
This highlights that the limitation of running on quantum hardware stems more from their current noise levels than from the ability of QAOA to perform effectively. We therefore expect that continued reduction in hardware noise, along with a combination of quantum error mitigation, suppression, and correction, will help overcome this limitation. In the current term though, we see that carrying out a noisy simulation can help inform the circuit depths that could be run on hardware before qubit noise becomes a significantly diminishing factor, as we have done here.
@@ II.3.6 Experiment Results
To execute our circuits on Garnet, we compiled and submitted our jobs through Infleqtion’s compilation software, Superstaq [21] which, in turn, submitted the job through AWS’s Braket. We ran QAOA for $p=3$ (based on the noisy simulations of Subsection II.3.5) and also $p=1$ and $p=5$ for comparison. Each QAOA circuit was executed with 20,000 shots (the maximum number of shots allowed by IQM for submissions to Garnet through AWS). The variational parameters $\alpha$ and $\beta$ were trained via classical simulation and the final quantum circuit was run on the quantum device using these optimized parameters. In the small case of this example experiment, classical computers can simulate the quantum circuit; however, at larger problem sizes, one may take advantage of parameter concentration (Section II.3.7) to train the variational parameters via classical simulation on a smaller, classically trainable, sub-problem and then run the quantum circuit for the full problem with these optimized parameters on a quantum computer. At a large enough scale, the quantum computer becomes necessary for running the final circuit as classical simulation of quantum circuits scales exponentially with the number of qubits.
Figure 12: Probabilities of the top 10 most probable states (rank ordered) measured on Garnet for $p=$ 1, 3, and 5. The opaque bars represent the optimal solution while the transparent bars represent non-optimal solutions.
Figure 12 shows the probabilities of the top ten most probable states (rank ordered) measured on Garnet for various $p$. The opaque bars represent the optimal solution while the translucent bars represent non-optimal solutions. Note that for each $p$, the most probable state measured was also the optimal state. The figure also demonstrates the two competing forces involved when selecting the best $p$. On the one hand, increasing $p$ increases the expressiveness of the circuit and therefore the effectiveness of QAOA. On the other hand, increasing $p$ also increases the depth of the quantum circuit, and therefore the noisiness of the results. For example, the plot for $p=1$ is sharp because of its short depth, yet low-probability because of its low expressivity. By contrast, the plot for $p=3$ is high probability because of its high expressivity, yet flat because of its long depth. The ‘sweet spot’ is at $p=3$ (as suggested by our noisy simulations from Subsection II.3.5), where the circuit is expressive enough to render high probabilities yet short enough to be sufficiently sharp. As one expands to larger, more expensive-to-run problem sizes, the importance of preemptively searching for the “sweet spot” for $p$ via noisy simulations of the circuit before running on hardware (Subsection II.3.5) becomes ever more apparent.
@@ II.3.7 Parameter Concentration
Figure 13: Full graph (all dots) and subgraph (red dots) used for the parameter concentration study.
Figure 14: Probability of measuring the best state versus number of layers $p$ for three instances: the subgraph solved on subgraph trained parameters $C_{\text{sub}}(\rho_{\text{sub}})$, the full graph solved on subgraph trained parameters $C_{\text{full}}(\rho_{\text{sub}})$, the full graph solved on full graph trained parameters $C_{\text{full}}(\rho_{\text{full}})$.
As noise often scales with the number of qubits/gates, and therefore problem size, it becomes harder to train QAOA as the problem size increases. This roadblock can be alleviated, however, if the problem exhibits parameter concentration [22], a phenomenon in which the QAOA parameters trained on a smaller instance of a problem perform well when applied to a larger instance of the same problem. Fortunately, we have numerical evidence suggesting parameter concentration for our particular problem of sphere packing. Figure 13 shows the full graph used for the hardware experiment. The dots colored red represent a subgraph which will serve as the “smaller problem instance” for our parameter concentration investigation. We train the parameters of both the subgraph QAOA circuit ($C_{\text{sub}}$) and the full graph QAOA circuit ($C_{\text{sub}}$), resulting in $\rho_{\text{sub}}$ and $\rho_{\text{full}}$, respectively. To test for parameter concentration, we compare the performance of the full graph circuit executed with the parameters trained on the subgraph $C_{\text{full}}(\rho_{\text{sub}})$ with that of the full graph circuit executed with the parameters trained on the full graph $C_{\text{full}}(\rho_{\text{full}})$. Figure 14 compares the performance of $C_{\text{full}}(\rho_{\text{sub}})$ and $C_{\text{full}}(\rho_{\text{full}})$ with respect to overlap with the optimal state for various numbers of layers $p$. It also includes the performance of the subgraph circuit executed with the parameters trained on the subgraph $C_{\text{sub}}(\rho_{\text{sub}})$. We can see that QAOA yields a larger overlap with the optimal state when solving the subgraph with the parameters trained on the subgraph $C_{\text{sub}}(\rho_{\text{sub}})$ when compared to solving the full graph with the parameters trained on the full graph $C_{\text{sub}}(\rho_{\text{sub}})$. This is expected as the subgraph has less parameters and is therefore easier to train. However, we can also see that QAOA on the full graph with parameters trained on the subgraph $C_{\text{full}}(\rho_{\text{sub}})$ performs only slightly worse than QAOA on the full graph with parameters trained on the full graph $C_{\text{full}}(\rho_{\text{full}})$, demonstrating parameter concentration. This suggests that QAOA may be able to solve real-world instances of this problem that are too large to train by executing QAOA on parameters trained on a trainable sub-instance.
@@ III Heterogeneous Packing
Using the encouraging results from real quantum hardware as a springboard, we generalize sphere packing to the heterogeneous case, for which good classical solutions are much harder to achieve. Consider the case of $s+1$ distinct radii $R=\{r_{n}\}_{n\in N}$ where $N=\{0,1,...,s\}$ and $r_{0}=0$ (a useful indicator for the nonexistence of a sphere). The state of the qubit (or set of qubits) at each node informs which sized sphere is centered at said node. To maximize the density of the packing, we minimize empty space subject to the constraint that the arrangement of spheres is non-overlapping. We formulate the cost Hamiltonian $H_{C}=H_{obj}+\lambda H_{con}$ for heterogeneous packing via two different formulations, which we call “First Quantization” and “Second Quantization” because of their analogies to the two levels of quantization in physics.
@@
$N_{vw}$ =$=\{n,m\in N_{v}\times N_{w}:|v-w|\leq r_{n}+r_{m}\}.$ (9)
Here, $N_{v}$ represents the index set of the set of radii that admit spheres of said radii to be placed at node $v$ without extending beyond the boundary of the cylinder. Similarly, $N_{vw}$ represents the index set of pairs of radii that would not admit the placement of spheres of said radii at nodes $v$ and $w$ without their overlapping. Here, $R_{b}$ is the radius of the cylinder. Converting these Hamiltonians to a qubit formulation (a requirement for any real hardware system today) requires only $\log_{2}|R|$ overhead via a binary encoding.
@@ IV Resource Scaling
Resource estimation for this problem is important to argue for the potential for quantum advantage over classical computation. To estimate this scaling of quantum resources for each problem formulation, we start by making several reasonable assumptions. Suppose we have a $d$-dimensional lattice with $q$ points per side placed within a $d$-dimensional cylinder with radius $R_{b}$. Furthermore, notate the set of radii, node sets, edge sets and the maximum radius as
@@
The number of CNOT gates to implement $e^{iH\gamma}$ in one QAOA layer is
$\sum_{r_{i},r_{j}\in R^{2}}|E_{r_{i}r_{j}}|$ $\leq|R|^{2}q^{d}\max_{r\in R}\left(\max_{v\in V_{r}}\deg(v)\right)$
$\leq|R|q^{d}\left(\frac{r_{m}q}{R_{b}}\right)^{d}$
@@
$R$ =$=\{r_{1},...,r_{s}\}$
$\Pi$ =$=\{1,...,\lceil\log_{2}\left(|R|+1\right)\rceil\}$
$V$ =$=\{v\}$
@@
where $E_{r_{i}r_{j}}$ has an edge $(v,w)$ if and only the distance between nodes $v$ and $w$ is less than $r_{i}+r_{j}$. With these definitions, the number of qubits is simply
$|\Pi||V|$ $\leq\left(\lceil\log_{2}\left(|R|+1\right)\rceil\right)q^{d}$
@@
$g(v,r_{i},w,r_{j})=\begin{cases}1&\text{if}\ vw\in E_{r_{i}r_{j}}\\ 0&\text{otherwise}\end{cases}$ (18)
@@
$\sum_{r_{i},r_{j}\in R^{2}}\sum_{v,w\in E_{r_{i}r_{j}}}$ $\leq|R|2^{2|\Pi|}\max_{r_{i},r_{j}\in R^{2}}\left(\max_{v\in V}\deg_{r_{i}r_{j}}(v)\right)$
$\leq|R|2^{\left({2\lceil\log_{2}\left(|R|+1\right\rceil}\right)}\left(\frac{r_{m}q}{R_{b}}\right)^{d}$
@@
Note that the bound on the single state penalty function is quite loose. Additionally, while the scaling may appear exponential in $d$, recall that for real-world applications $d$ is only $2$ or $3$, keeping the quantum resource scaling for both formulations bounded by a low order polynomial in all its components. Note that compared to the “first-quantization” formulation, the “second-quantization” formulation trades an increase in the CNOT scaling degree for an exponential decrease in the number of qubits. The best formulation to use is therefore dependent on which resource will incur a larger noise penalty for near-term quantum devices, their number of qubits or number of gates. Either way, we have shown that the quantum resources for this problem only scale polynomially with increased problem size - in contrast to classical computers, whose complexity grows exponentially as this problem is NP-hard.